OpenStudy (hy123):

For the real-valued functions f(x)=sqrt(5x+20) and g(x)=x-2 , find the composition f o g(x) and specify its domain using interval notation.

OpenStudy (hy123):

\[f(x)=\sqrt{5x+20}\] g(x)= x-2

OpenStudy (hy123):

Hey gane

OpenStudy (hy123):

\[\sqrt{5(x-2)+20}\]

ganeshie8 (ganeshie8):

yep.. this is f(g(x)) we need to find its domain ?

OpenStudy (hy123):

\[\sqrt{5x+10}\]?

OpenStudy (hy123):

Do i need to factor or just leave it as it is? And domain would be (-infinity,-2) union (-2,+infinity)?

ganeshie8 (ganeshie8):

check again..

ganeshie8 (ganeshie8):

domain of a function is the set of all values we can supply as input..

ganeshie8 (ganeshie8):

since its a real valued function, 5x+10 >= 0

OpenStudy (hy123):

Oh wait... Uhh.... (-2,+infinity)?

ganeshie8 (ganeshie8):

its [-2, infiniti) -2 included... sqrt(0) is 0. its defined !!

OpenStudy (hy123):

Arg -.-